Show that if A is an n × n matrix, then AAT and A + AT are symmetric


1
Expert's answer
2021-05-13T03:13:48-0400

"Solution:
\\ A ~is ~an~ n × n~ matrix~i.e. ~square~ matrix. 
\\If~ A^T=A ~then ~matrix ~ A~ is ~ symmetric.
\\Let~K=AA^T
\\\therefore K^T=(AA^T)^T
\\~~~~~~~~~~~=(A^T)^TA^T~~~~~[Since~(AB)^T=B^TA^T]
\\~~~~~~~~~~~=AA^T~~~~[Since ~(A^T)^T=A]
\\~~~~K^T=K
\\Hence ~ AA^T~is ~symmetric.
\\Now ~let~us ~ consider ~C=A+A^T
\\\therefore C^T=(A+A^T)^T
\\~~~~~~~~~~~=A^T+(A^T)^T
\\~~~~~~~~~~~=A^T+A~~~~~~~[Since~(A^T)^T=A ]
\\~~~~~~~~~~~=A+A^T ~~~~~~~[A+A^T=A^T+A~~~Commutative Property]
\\~~~~C^T=C
\\Hence~ A+A^T~ is~ symmetric.
\\Hence~if ~A~ is ~an~ n × n ~matrix,\\ then ~AA^T ~and~ A + A^T ~are~ symmetric."